# The Evolution and Internal Structure of Neptunes and Sub-Neptunes

### v1

@@ 4 Discussion
While our study presents a step forward towards the understanding of Neptunes and sub-Neptunes, some simplifying assumptions have to be made as we discuss below. First, we note that the planetary cooling strongly depends on the atmospheric model. In this work we used a standard semi-gray irradiated atmosphere model (Guillot & Havel, 2011). Our choice of atmospheric model represents a compromise between accuracy and complexity. The semi-gray approximation has been shown to deviate by no more than $\lesssim 10\%$ from full numerical solutions to the radiative transfer problem across a wide range of gravities and effective temperatures (Parmentier et al., 2015). However, a different model would change the inferred thermal energy flux through the atmosphere, leading to a different temperature at the bottom of the atmosphere. If this flux is significantly different compared to the one we calculate here, the radius evolution would also be affected. Since the contraction of the outer envelope depends on the energy balance between the flux coming from the deep interior and the flux that is emitted through the atmosphere, our results are still valid in terms of the relative differences between different model assumptions.
@@
Third, for simplicity, we kept the composition profile fixed over time and only consider the effect of the energy transport from convection. In reality, convective mixing erodes composition gradients (e.g. Knierim & Helled, 2024; Tejada Arevalo, 2025). Including the effect of convective mixing would affect the inferred evolution, especially for the composition profile with a wide transition region. Such a composition gradient could mix from outside inward, and the composition may start building a step-like profile with narrow non-convective regions separating convective regions. As already stated in the introduction, Uranus is suspected to have a stable non-convective boundary layer not too far inside its interior. Furthermore, if planet formation happens relatively cold (low entropy), mixing is likely to be less efficient, and stable composition gradients can be expected in many planets. Therefore, as long as there is a stable non-convective region, it is clear that the choice of the conductivity is significant and should be carefully considered. We hope to further improve on that in a future study that self-consistently treats convection and mixing.
@@
Fourth, due to lack of data, the assumed material of the EoS is not always the same as the material in our conductivity models. We took four choices for the conductivity models that all assume a different material. This material is different from the water-rock mixture we assume for the EoS and does not scale with the metallicity. This inconsistency is hard to improve as the data for the conductivities are spars. We hope that further investigations of the thermal conductivity of different materials will provide the required information, allowing for more consistent planetary models.
@@
Fifth, although we investigated multiple cases with different masses, the initial temperature profile in the models is nearly adiabatic. However, formation models suggest that composition gradients form during the planetary formation process (Valletta & Helled, 2020, 2022). Therefore, Neptunes and sub-Neptunes could have temperature profiles that differ from the adiabatic ones after their formation. The composition profile we used here is based on the idea of having a smooth transition between the central metallicity and an envelope metallicity similar to what is expected from Uranus interior models. Because mixing likely starts from outside in, we assume a flat composition profile for the envelope, which might have been an eroded composition gradient. We suggest that future studies include large ranges of bulk metallicities and internal structures.
@@
Sixth, the planetary evolution depends on the choice of the EoS. In Appendix G, we compare the radius evolution obtained with different H–He EoS and find that simulations using the SCvH EoS (Saumon et al., 1995) yield radii up to $\lesssim 2\%$ larger after 10 Gyr compared to those with the CMS EoS. Variations in the EoS of heavy elements would also affect the results. For example, Howard et al. (2025) compared two water EoS for static adiabatic models for Uranus, which leads to a temperature difference of $\sim$1000K at a few hundred kbar. Given the large uncertainty in the planetary composition and the unknown ratio between different elements (e.g., ice and rock), the uncertainty of the EOS is expected to be small. At the same time, we note that a systematic study of the uncertainties arising from both the EoS and the assumed materials is desirable.
@@
Seventh, for simplicity, we do not consider demixing and settling, which are processes that could take place and affect the evolution of Sub-Neptunes and Neptunes. For example, the phase separation of water (Bailey & Stevenson, 2021; Cano Amoros et al., 2024; Howard et al., 2025) or carbon-bearing species (He et al., 2022; Cheng et al., 2023; Militzer, 2024) could delay the planetary cooling and lead to radius inflation. We also note that progress in characterizing material properties of various mixtures in planetary conditions is required. We hope to address these topics in future research.
@@
Finally, we note that for planets with a significant amount of water in the deep interior, we used the easy-to-use fits for the conductivity from French & Redmer (2017) for the electronic contribution and French (2019) for the vibrational contribution. Scheibe et al. (2021) employ the same models for thermal conductivity and conclude that the presence of additional materials would enhance the conductivity relative to pure water. Thus, using the conductivity of pure-water provides a lower bound. For example, the higher electrical conductivity of ammonia compared to water could lead to higher thermal conductivity (Ravasio et al., 2021). It is therefore clear that in order to perform reliable planetary evolution simulations, one should use the appropriate conductivity data. We suggest that it is highly desriable to investigate the properties of different H–C–N–O mixtures at planetary conditions.
Our work shows the similarly to the equation of state, the material conductivity plays an important role in planetary evolution and therefore also for planetary characterization.
@@ 5 Summary and conclusions
Our galaxy is full with Neptunes and sub-Neptunes but their characterization remains a challenge. Often, these planets are modeled assuming a differentiated structure and adiabatic interior. In this work, we investigated how the planetary thermal evolution, and therefore also the inferred internal structure, is affected when the planets consist of non-convective layers.
@@ Acknowledgements.
Figure 10: Radiative opacity as a function of temperature for constant $\log R$. The left, middle and right panels correspond to metallicities of Z=0.02, 0.20Z=0.02,\,0.20, and 0.95, respectively. The dashed line shows the radiative opacity from Freedman et al. (2014) as included in MESA and the solid line shows the custom-made AESOPUS2.1 tables using the web interface (Marigo et al. 2024). The three different colors show different values for $\log R$. Note that the AESOPUS2.1 tables are only available up to $\log R=6$, resulting in the same AESOPUS2.1 opacities for $\log R\geq 6$ for a given temperature (orange and green solid-line).
@@ Appendix B Vibrational conductivity
Ross et al. (1984) reviewed the phonon model for a completely anharmonic crystal and derived the temperature and density scaling of the thermal conductivity of liquids under high pressure. We follow the formulation of Stamenković et al. (2011), such that the partial derivatives of the vibrational conductivity $k_{\text{vib}}$ are given by:
@@
with $\gamma=\left(\partial\ln\nu/\partial\ln\rho\right)_{T}$ being the Grüneisen parameter for the average vibrational frequency $\nu$. The parameter $\gamma$ can be approximated using the thermodynamic Grüneisen parameter:
@@
where $\alpha=-\left(\partial\ln\rho/\partial\ln T\right)_{P}$ is the thermal expansivity, $K_{T}=-\left(\partial\ln\rho/\partial\ln P\right)_{T}$ the isothermal compressibility, and $c_{V}$ the specific heat capacity at constant volume. Given a reference conductivity $k_{\text{vib, ref}}(T_{\text{ref}},\rho_{\text{ref}})$ and an EOS to calculate $\gamma_{\text{TH}}$, one can construct the vibrational conductivity from:
@@ 2. Step 2:
    Next, we relax the composition to $Z_{\text{bulk}}=0.94$ using relax_initial_Z to avoid convergence issues for planets with masses of $M_{\text{p}}\lesssim 8$ M. Ideally, one would set the metallicity in this step to the final central metallicity to ensure a consistent entropy, which is set in step Step 4. However, this led to issues, which is why we set $Z_{\text{bulk}}$ to a lower value. This is accounted for by setting a slightly higher entropy in step Step 4. The model is forced to follow the adiabat.
@@ 3. Step 3:
    Now we use the relax option relax_initial_mass to shrink the current model down to its desired mass $M_{\text{p, i}}$. This step can be difficult if $M_{\text{p, i}}$ is relatively small. In this case either increasing the metallicity as described in Step 2 can help or changing the energy state by setting a different flat entropy profile before this step as described Step 4. Again, we force the model to follow the adiabat.
@@ 4. Step 4:
    We set the initial energy budget, which we parametrize by the central specific entropy $s_{\text{center, i}}$. Using the option relax_initial_entropy MESA applies an extra specific heat $Q_{l}$ in each layer $l$ according to:
@@
    with $s_{l}$ being the current specific entropy, $E_{l}$ the current specific energy and $\tau_{\text{relax entropy}}=1\times 10^{-3}$ yr an relax timescale. Usually, the adiabat i.e. a constant entropy profile, agrees with the profile that is predicted by the mixing length theory. But especially for the outer parts the entropy profile can significantly deviate from an adiabat. This can lead to convergence issues where the solver reacts with rapidly changing profiles. We avoid this issue by forcing the model to follow the adiabatic gradient given by the EoS. The specific entropy depends on the metallicity, and our final composition profile features a central metallicity of $Z_{\text{int, i}}=1$, which differs from the value set in step Step 2. As a result, in this step we assign a higher entropy such that the initial model after relaxing the composition has the desired entropy.
@@ 5. Step 5:
    We consider two cases for the composition profile, one with a wide transition region (transition over $\Delta q=0.1$) and one with a narrow transition region (transition over $\Delta q=0.01$). Both profiles are fixed to the same bulk metallicity of $Z_{\text{bulk}}=0.95$ with an envelope metallicity of $Z_{\text{env}}=0.2$.
@@ Appendix G The effect of the H-He EoS
In this study we used the CMS EoS with non-ideal interactions for H-He (see Section 2 for details). Here, we present the planetary evolution when using the SCvH EoS for H-He (Saumon et al. 1995) instead of CMS. Figure 15 shows the radius evolution for the models with a warm start and a wide composition gradient using both EoS. We find small deviations of $\lesssim 2\%$ for the planetary radius after 10 Gyr. Comparing the water EoS using MESA is currently not possible due to the rather complex implementation procedure (e.g., Helled et al., 2025, submitted).

### v2

@@ 4 Discussion
While our study presents a step forward toward the understanding of Neptunes and sub-Neptunes, some simplifying assumptions have to be made, which we discuss below. First, we note that the planetary cooling strongly depends on the atmospheric model. In this work we used a standard semi-gray irradiated atmosphere model (Guillot & Havel, 2011). Our choice of atmospheric model represents a compromise between accuracy and complexity. The semi-gray approximation has been shown to deviate by $\lesssim 10\%$ from full numerical solutions to the radiative transfer problem across a wide range of gravities and effective temperatures (Parmentier et al., 2015). However, a different model would change the inferred thermal energy flux through the atmosphere, leading to a different temperature at the bottom of the atmosphere. If this flux is significantly different compared to the one we calculate here, the radius evolution would also be affected. Since the contraction of the outer envelope depends on the energy balance between the flux coming from the deep interior and the flux that is emitted through the atmosphere, our results are still valid in terms of the relative differences between different model assumptions.
@@
Third, for simplicity, we kept the composition profile fixed over time and only considered the effect of the energy transport from convection. In reality, convective mixing erodes composition gradients (e.g. Knierim & Helled, 2024; Tejada Arevalo, 2025). Including the effect of convective mixing would affect the inferred evolution, especially for the composition profile with a wide transition region. Such a composition gradient could mix from outside inward, and the composition may start building a step-like profile with narrow non-convective regions separating convective regions. As already stated in the introduction, Uranus is suspected to have a stable non-convective boundary layer not too far inside its interior. Furthermore, if a planet forms cold (low entropy), mixing is likely to be less efficient, and stable composition gradients can be expected in many planets. Therefore, as long as there is a stable non-convective region, it is clear that the choice of the conductivity is significant and should be carefully considered. We hope to further improve on this choice in a future study that self-consistently treats convection and mixing.
@@
Fourth, due to a lack of data, the assumed material of the EoS is not always the same as the material in our conductivity models. We took four choices for the conductivity models that all assume a different material. This material is different from the water-rock mixture we assumed for the EoS and does not scale with the metallicity. This inconsistency is hard to improve, as the data for the conductivities are sparse. We hope that further investigations of the thermal conductivity of different materials will provide the required information, allowing for more consistent planetary models.
@@
Fifth, although we have investigated multiple cases with different masses, the initial temperature profile in the models is nearly adiabatic. However, formation models suggest that composition gradients form during the planetary formation process (Valletta & Helled, 2020, 2022). Therefore, Neptunes and sub-Neptunes could have temperature profiles that differ from the adiabatic ones after their formation. The composition profile we used here is based on the idea of having a smooth transition between the central metallicity and an envelope metallicity similar to what is expected from Uranus interior models. Because mixing likely starts from the outside in, we assumed a flat composition profile for the envelope, which might have been an eroded composition gradient. We suggest that future studies include large ranges of bulk metallicities and internal structures.
@@
Sixth, the planetary evolution depends on the choice of the EoS. In Appendix G, we compare the radius evolution obtained with different H–He EoSs and find that simulations using the SCvH EoS (Saumon et al., 1995) yield radii up to $\lesssim 2\%$ larger after 10 Gyr compared to those with the CMS EoS. Variations in the EoS of heavy elements would also affect the results. For example, Howard et al. (2025) compared two water EoSs for static adiabatic models for Uranus, which leads to a temperature difference of $\sim$1000K at a few hundred kbar. Given the large uncertainty in the planetary composition and the unknown ratio between different elements (e.g., ice and rock), the uncertainty of the EOS is expected to be small. At the same time, we note that a systematic study of the uncertainties arising from both the EoS and the assumed materials is desirable.
@@
Seventh, for simplicity, we did not consider demixing and settling, which are processes that could take place and affect the evolution of sub-Neptunes and Neptunes. For example, the phase separation of water (Bailey & Stevenson, 2021; Cano Amoros et al., 2024; Howard et al., 2025) or carbon-bearing species (He et al., 2022; Cheng et al., 2023; Militzer, 2024) could delay the planetary cooling and lead to radius inflation. We also note that progress in characterizing material properties of various mixtures in planetary conditions is required. We hope to address these topics in future research.
@@
Finally, we note that for planets with a significant amount of water in the deep interior, we used the easy-to-use fits for the conductivity from French & Redmer (2017) for the electronic contribution and those of French (2019) for the vibrational contribution. Scheibe et al. (2021) employed the same models for thermal conductivity and conclude that the presence of additional materials would enhance the conductivity relative to pure water. Thus, using the conductivity of pure water provides a lower bound. For example, the higher electrical conductivity of ammonia compared to water could lead to higher thermal conductivity (Ravasio et al., 2021). It is therefore clear that in order to perform reliable planetary evolution simulations, one should use the appropriate conductivity data. We suggest that it is highly desirable to investigate the properties of different H–C–N–O mixtures at planetary conditions. Our work shows that, similar to the EoS, the conductivity of material plays an important role in planetary evolution and therefore also in planetary characterization.
@@ 5 Summary and conclusions
Our Galaxy contains many Neptunes and sub-Neptunes, but their characterization remains a challenge. Often, these planets are modeled assuming a differentiated structure and adiabatic interior. In this work, we have investigated how the planetary thermal evolution, and therefore also the inferred internal structure, is affected when the planets consist of non-convective layers.
@@ Acknowledgements.
Figure 10: Radiative opacity as a function of temperature for constant $\log R$. The left, middle and right panels correspond to metallicities of Z=0.02, 0.20Z=0.02,\,0.20, and 0.95, respectively. The dashed line shows the radiative opacity from Freedman et al. (2014) as included in MESA and the solid line shows the custom-made AESOPUS2.1 tables using the web interface (Marigo et al. 2024). The three different colors show different values for $\log R$. We note that the AESOPUS2.1 tables are only available up to $\log R=6$, resulting in the same AESOPUS2.1 opacities for $\log R\geq 6$ for a given temperature (orange and green solid-line).
@@ Appendix B Vibrational conductivity
Ross et al. (1984) reviewed the phonon model for a completely anharmonic crystal and derived the temperature and density scaling of the thermal conductivity of liquids under high pressure. We follow the formulation of Stamenković et al. (2011), such that the partial derivatives of the vibrational conductivity $k_{\text{vib}}$ are given by
@@
with $\gamma=\left(\partial\ln\nu/\partial\ln\rho\right)_{T}$ being the Grüneisen parameter for the average vibrational frequency $\nu$. The parameter $\gamma$ can be approximated using the thermodynamic Grüneisen parameter
@@
where $\alpha=-\left(\partial\ln\rho/\partial\ln T\right)_{P}$ is the thermal expansivity, $K_{T}=-\left(\partial\ln\rho/\partial\ln P\right)_{T}$ the isothermal compressibility, and $c_{V}$ the specific heat capacity at constant volume. Given a reference conductivity $k_{\text{vib, ref}}(T_{\text{ref}},\rho_{\text{ref}})$ and an EOS to calculate $\gamma_{\text{TH}}$, one can construct the vibrational conductivity from
@@ 2. Step 2:
    Next, we relax the composition to $Z_{\text{bulk}}=0.94$ using relax_initial_Z to avoid convergence issues for planets with masses of $M_{\text{p}}\lesssim 8$ M. Ideally, one would set the metallicity in this step to the final central metallicity to ensure a consistent entropy, which is set in Step 4. However, this led to issues, which is why we set $Z_{\text{bulk}}$ to a lower value. This is accounted for by setting a slightly higher entropy in Step 4. The model is forced to follow the adiabat.
@@ 3. Step 3:
    Now we use the relax option relax_initial_mass to shrink the current model down to its desired mass $M_{\text{p, i}}$. This step can be difficult if $M_{\text{p, i}}$ is low. In this case, increasing the metallicity as described in Step 2 can help. Or before changing the mass one can set a different entropy profile as described in Step 4. Again, we force the model to follow the adiabat.
@@ 4. Step 4:
    We set the initial energy budget, which we parametrize by the central specific entropy $s_{\text{center, i}}$. Using the option relax_initial_entropy MESA applies an extra specific heat $Q_{l}$ in each layer $l$ according to
@@
    with $s_{l}$ being the current specific entropy, $E_{l}$ the current specific energy and $\tau_{\text{relax entropy}}=1\times 10^{-3}$ yr an relax timescale. Usually, the adiabat, i.e., a constant entropy profile, agrees with the profile that is predicted by the mixing length theory. But especially for the outer parts the entropy profile can significantly deviate from an adiabat. This can lead to convergence issues where the solver reacts with rapidly changing profiles. We avoid this issue by forcing the model to follow the adiabatic gradient given by the EoS. The specific entropy depends on the metallicity, and our final composition profile features a central metallicity of $Z_{\text{int, i}}=1$, which differs from the value set in step Step 2. As a result, in this step we assign a higher entropy such that the initial model after relaxing the composition has the desired entropy.
@@ 5. Step 5:
    We consider two cases for the composition profile, one with a wide transition region (transition over $\Delta q=0.1$) and one with a narrow transition region (transition over $\Delta q=0.001$ [^2] The published version contained a typo, stating the width of the narrow gradient as $\Delta q=0.01$. The correct value used in the simulations is $\Delta q=0.001$.). Both profiles are fixed to the same bulk metallicity of $Z_{\text{bulk}}=0.95$ with an envelope metallicity of $Z_{\text{env}}=0.2$.
@@ Appendix G The effect of the H-He EoS
In this study we used the CMS EoS with non-ideal interactions for H-He (see Section 2 for details). Here, we present the planetary evolution when using the SCvH EoS for H-He (Saumon et al. 1995) instead of CMS. Figure 15 shows the radius evolution for the models with a warm start and a wide composition gradient using both EoS. We find small deviations of $\lesssim 2\%$ for the planetary radius after 10 Gyr. Comparing the water EoS using MESA is currently not possible due to the rather complex implementation procedure (e.g. Helled et al. 2025).